OpenStudy (anonymous):

Evaluate 3ab for a = 2 and b = 3. 6 9 18

OpenStudy (anonymous):

haaaaaiii

OpenStudy (anonymous):

3*2*3

OpenStudy (anonymous):

3*3=6

OpenStudy (anonymous):

3*2=6

OpenStudy (anonymous):

6*3=18

OpenStudy (anonymous):

it would be 18

OpenStudy (anonymous):

ty